One more than twice the square of a number is 73. Translate to an equation and solve
Answer by checkley79(3341)

2x^2+1=73
2x^2=73-1
2x^2=72
x^2=72/2
x^2=36
x=sqrt36
x=6 ans. for the number.
Proof:
2*6^2+1=73
2*36+1=73
72+1=73
73=73